OpenStudy (anonymous):

Compute the curvature and the principal unit normal of the elliptical helix described by the equations: x=cost, y=2sint, z=t

OpenStudy (anonymous):

i differentiated each component and got: x'=-sint, y'=2cost, and z=1

OpenStudy (anonymous):

i belive that should be my v(t), now i need to find the magniutude so i computed: \[\sqrt{\sin^2t+4cost+1}\]

OpenStudy (anonymous):

how do i go about simplifying that

OpenStudy (anonymous):

im sorry, it should be 4cos^2(t)

OpenStudy (dumbcow):

not sure you can really simplify that cos^2 = 1-sin^2

OpenStudy (anonymous):

is the right process i am doing, to compute the k, curvature

OpenStudy (dumbcow):

or sin^2 = 1-cos^2 i believe so but i would have to look it up , i don't work with these very often sorry

OpenStudy (anonymous):

alright

OpenStudy (dumbcow):

is the curvature pretty much finding the gradient

OpenStudy (anonymous):

i belive it is v/magnitude of v

OpenStudy (dumbcow):

i found this which may help, looks like you need to take 2nd derivatives http://en.wikipedia.org/wiki/Curvature#One_dimension_in_three_dimensions:_Curvature_of_space_curves